# Pull-Based Query Scheduling for Goal-Oriented Semantic Communication

### v1

@@ V-B1 Computing $\pi$
To derive the scheduling policies, we adapt two prominent on-policy DRL algorithms, advantage actor-critic (A2C) [35] and proximal policy optimization (PPO) [36], along with an off-policy algorithm, Deep Q-network (DQN) [37]. These algorithms are particularly well-suited for decision-making in high-dimensional spaces.
In each algorithm, the Q-function or the corresponding value function is computed based on the net reward defined in (V-A).
@@ V-B2 Computing $\mu^{*}$
We employ the same method as in Algorithm 1 to determine the optimal Lagrange multiplier for a given policy from the inner loop. In this context, the bisection search method is employed in the outer loop to iteratively and gradually identify the minimum Lagrange multiplier that meets the cost constraint of the scheduling problem $\mathcal{P}$, as formulated in (III-A).
@@ VI-A Setup and Assumptions
Number of slots $T$ $10^{3}$ Parameters shaping $v_{\rm cpt}(\cdot)$ $\alpha_{\rm cpt}$ $0.5$
Number of SAs $N$ $4$ $\beta_{\rm cpt}$
Number of attributes $M$ $2$ $\lambda_{\rm cpt}$ $2$
Number of AAs $K$ $4$ Refs. point $\operatorname{GoE}_{\rm ref}$ $0.2$
Corr. observation probability $p_{{\rm{o}},nm},$ $\forall n,m$ $0.8$ Cost per query $f_{c}(1)$ $0.5$
Required attributes $\lvert\mathcal{M}_{k}\rvert,\forall k$ $2$ Cost flex. index $C_{\rm flex}$ $0.75$
Erasure probability $p_{{\rm{e}},n},\forall n$ $0.2$ Maximum AoI $\Delta_{\rm max}$ $4$
@@
Shape parameters for $g_{m}(\cdot;\cdot),\forall m$$\{\alpha_{1},\alpha_{2}\}$$\{0.5,2\}$ Tolerance sens. $\epsilon_{\mu}$
@@
Size of usefulness $\lvert\mathcal{U}\rvert$ $4$
To implement the DRL algorithms, namely A2C, PPO, and DQN, we follow the default hyperparameter settings as considered in their respective original papers, i.e., [35], [36], and [37]. However, we adjust the discount factor according to Table I. We employ the Adaptive Moment Estimation (Adam) optimizer for both DQN and PPO, whereas the Root Mean Square Propagation (RMSprop) optimizer is used for A2C. For each DRL algorithm, a model is trained over $100$ episodes of interaction. Each episode comprises $10,000$ time slots, generating a total of $T_{\rm e}=10^{6}$ environment steps. The key hyperparameters for these algorithms, along with their initial configurations, are summarized in Table II.
@@
Learning rate $10^{-4}$ $7\!\times\!10^{-4}$ $3\!\times\!10^{-4}$
@@
Rollout buffer size $5\!\times\!8$ $1\!\times\!2048$
@@
Bias-variance trade-off $1$ $0.95$
@@ VI-B Results and Discussion
Expanding on the time-variant effectiveness provided by different approaches, Fig. 4 depicts the instantaneous CPT-based total GoE at each slot over the interval $[0,50]$. It highlights that, unlike other approaches that show significant fluctuations in performance, both model-based and model-free effect-aware scheduling demonstrate consistent and stable performance, ensuring a minimum level of effectiveness. Specifically, in Fig. 4 (b), effect-aware scheduling consistently maintains a CPT-based total GoE of $0.29$ or higher throughout the observed period after passing the initial state, i.e., $t=0$. In contrast, using LWGF, WRR, uniform, and Markovian, the CPT-based total GoE could decrease to as low as $0.23$, $0.17$, $-0.35$, and $-0.35$, respectively, as shown in Fig. 4 (a). This confirms that effect-aware scheduling outperforms the benchmark approaches in consistently maintaining reliable and stable performance.
@@
To study the interplay between the average CPT-based total GoE and its corresponding reference point, i.e., $\operatorname{GoE}_{\rm ref}$, we plot Fig. 5. This graph depicts the effect of increasing the reference point on the effectiveness and performance of different scheduling approaches. As $\operatorname{GoE}_{\rm ref}$ increases, the effectiveness delivered by all approaches decreases. This decrease can be attributed to the activation of the loss component in the CPT-based value function at higher levels of total GoE. However, as the reference point increases, the performance gap between effect-aware scheduling and the other approaches widens, reaching its peak at $\operatorname{GoE}_{\rm ref}=0.5$. Beyond this critical point, the performance of both model-based and model-free approaches starts to degrade. Notably, after $\operatorname{GoE}_{\rm ref}=0.95$, the benchmark approaches outperform all effect-aware ones. This reversal may result from the increasing reference point reducing the difference between the net rewards across $\lvert\mathcal{S}\rvert=256$ states, which could distort the optimal value of the Lagrange multiplier. Consequently, this misguides the scheduling policies, whether model-based or model-free. As a result, effect-aware scheduling performs poorly as the reference point approaches the upper bound.
@@
Regarding the scalability of model-free effect-aware scheduling with a large number of attributes, as depicted in Fig. 7, we analyze how increasing the query limit affects effectiveness. Under the constraint $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 1$ in (III-A), the hub is essentially restricted to sending at most one query during each sampling interval $t$. Fig. 8 illustrates the average CPT-based total GoE over $1,000$ time slots as a function of the query limit, using model-free approaches. Their performance is compared against that of LWGF and uniform scheduling. [^2] Setting the query limit to more than one increases the complexity of the model-free solution and reduces its appeal for exploration in this context, as this requires corresponding modifications to the CMDP model in Section III-B. To plot the corresponding curves, we assume $\lvert\mathcal{M}_{k}\rvert=3,\forall k$ and vary the query limit from $1$ to $3$. [^3] Given that update packets are typically small, we assume that multiple attribute updates can be transmitted simultaneously through orthogonal channels or using time-division techniques. The figure shows that increasing the query limit enhances effectiveness across all approaches. Notably, model-free effect-aware scheduling consistently outperforms LWGF on average, particularly when $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 2,\forall t$. Both approaches demonstrate a significant performance advantage over uniform scheduling. When $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 3$, the hub can query all attributes simultaneously while still adhering to the cost constraint. At this point, the performance of model-free effect-aware scheduling and LWGF converge, becoming nearly identical.
@@ VII Conclusion
In this paper, we studied the query scheduling problem in end-to-end status update systems operating under a pull-based model. We considered multiple SAs observing a source with various attributes and communicating their observations in the form of updates in response to queries from the hub. The hub is responsible for deciding when to query updates and for constructing a knowledge base to enable AAs to achieve their heterogeneous goals. We introduced the GoE metric and formulated the scheduling problem based on this metric, integrating a risk-averse framework. Specifically, we aimed to maximize the expected discounted sum of the CPT-based total GoE while ensuring compliance with a cost constraint. To solve this problem and design effect-aware scheduling policies, we developed model-based and model-free solutions, supported by their respective iterative algorithms. We then evaluated the performance of these solutions in terms of effectiveness and compared them against benchmark methods. Our results demonstrated that the model-based method could improve the efficiency of sent queries by up to $20\%$ while preserving a level of effectiveness comparable to the best benchmark method, namely LWGF. Conversely, model-free solutions showed the potential to enhance long-term effectiveness by up to $9.2\%$ than the other approaches. This improvement was particularly pronounced under strict cost constraints, where effect-aware scheduling policies can significantly boost effectiveness and outperform others. Our findings highlighted that the model-free solutions excel in scalability while maintaining high levels of effectiveness, making them particularly appealing for large-scale or dynamic decision-making systems.

### v2

@@ V-B1 Computing $\pi$
To derive scheduling policies, we adapt Deep Q-Network (DQN) [36], Advantage Actor–Critic (A2C) [37], and Proximal Policy Optimization (PPO) [38]. We begin with the off-policy DQN algorithm, which approximates the optimal state–action value function $Q^{*}_{\pi}(s(t),a(t))$ using a deep neural network. The Q-values are updated by minimizing the temporal-difference loss between predicted and target Q-values, as defined by the Bellman equation. In our formulation, the Q-function and value function are computed as in (11)–(12), based on the net reward defined in (V-A). To improve training stability, DQN incorporates two key techniques: experience replay, which mitigates sample correlation by learning from a buffer of past transitions, and a target network, which is updated at a slower rate to stabilize Q-value estimates.
In contrast, A2C is an on-policy, synchronous actor–critic method in which the actor learns a stochastic policy $\pi(s(t))$, and the critic estimates the value function $V_{\pi}(s(t))$ for $s(t)\in\mathcal{S}$. To reduce the variance of policy gradient updates, it employs the advantage function $A_{\pi}(s(t),a(t))=Q_{\pi}(s(t),a(t))-V_{\pi}(s(t))$, which measures how much better action $a(t)$ is relative to the expected value of state $s(t)$ at the $t$-th time slot. A2C performs multiple parallel rollouts across environments, aggregates gradients, and synchronously updates both the actor and critic networks. This design fosters more stable training while preserving the sample efficiency of actor–critic methods.
Finally, PPO, another on-policy algorithm, also follows an actor–critic architecture while enhancing training stability through constrained policy updates. It directly optimizes a stochastic policy by maximizing a clipped surrogate objective, which penalizes large deviations from the previous policy. This clipping mechanism limits abrupt changes in the action probability ratio between successive policies, thereby mitigating the risk of overly aggressive updates that can destabilize learning. By enforcing conservative yet effective policy adjustments, PPO improves upon earlier policy gradient methods and can offer more reliable convergence across a wide range of tasks.
These algorithms exhibit complementary strengths and trade-offs: A2C and PPO, due to their on-policy design, provide stable policy updates in dynamic environments, making them effective in scenarios where adaptability is crucial. However, this comes at the cost of lower sample efficiency, as they require extensive interaction with the environment, potentially slowing convergence. In contrast, DQN benefits from off-policy learning and experience replay, enabling higher sample efficiency and making it particularly well-suited for discrete action spaces. Nevertheless, DQN may face challenges in high-dimensional or partially observable environments, where its value-based estimation can lead to instability and limited generalization across diverse state–action pairs. [^3] More recent or specialized algorithms, such as Soft Actor–Critic (SAC) and Twin Delayed Deep Deterministic Policy Gradient (TD3), are excluded from consideration, as they are tailored for continuous control tasks and introduce additional complexity without offering clear benefits for the discrete scheduling objective addressed in this work.
@@ V-B2 Computing $\mu^{*}$
We employ the same approach as in Algorithm 1 to determine the optimal Lagrange multiplier for a given policy from the inner loop. In this context, the bisection search method is employed in the outer loop to iteratively and gradually identify the minimum Lagrange multiplier that meets the cost constraint of the scheduling problem $\mathcal{P}$, as formulated in (III-A).
@@ VI-A Setup and Assumptions
Number of slots $T$ $10^{3}$
Parameters
shaping $v_{\rm cpt}(\cdot)$
$\alpha_{\rm cpt}$ $0.5$
Number of SAs NN 44 $\beta_{\rm cpt}$
Number of attributes $M$ 22 $\lambda_{\rm cpt}$ 22
Number of AAs $K$ 44 Refs. point $\operatorname{GoE}_{\rm ref}$ $0.2$
Corr. observation
probability
$p_{{\rm{o}},nm},$
$\forall n,m$
$0.8$ Cost per query $f_{c}(1)$ $0.5$
Required attributes $\lvert\mathcal{M}_{k}\rvert,\forall k$ 22 Cost flex. index $C_{\rm flex}$ $0.75$
Erasure probability $p_{{\rm{e}},n},\forall n$ $0.2$ Maximum AoI $\Delta_{\rm max}$ 44
@@
Shape parameters
for $g_{m}(\cdot;\cdot),\forall m$
$\{\alpha_{1},\alpha_{2}\}$ $\{0.5,2\}$ Tolerance sens. $\epsilon_{\mu}$
@@
Size of usefulness $\lvert\mathcal{U}\rvert$ 44
To implement the DRL algorithms, namely A2C, PPO, and DQN, we mostly follow the default hyperparameter settings as considered in their respective original papers, i.e., [37], [38], and [36]. However, we adjust the discount factor according to Table I. We employ the Adaptive Moment Estimation (Adam) optimizer for both DQN and PPO, whereas the Root Mean Square Propagation (RMSprop) optimizer is used for A2C. For each DRL algorithm, a model is trained over $100$ episodes of interaction. Each episode comprises $10,000$ time slots, generating a total of $T_{\rm e}=10^{6}$ environment steps. The key hyperparameters for these algorithms, along with their initial configurations, are summarized in Table II.
@@
Learning rate $10^{-4}$ $3\!\times\!10^{-4}$ $3\!\times\!10^{-4}$
@@
Rollout buffer size $20\!\times\!8$ $1\!\times\!2048$
@@
Bias-variance trade-off 11 $0.95$
@@ VI-B Results and Discussion
Expanding on the time-variant effectiveness provided by different approaches, Fig. 4 depicts the instantaneous CPT-based total GoE at each slot over the interval $[0,50]$. It highlights that, unlike other approaches that show significant fluctuations in performance, both model-based and model-free effect-aware scheduling demonstrate consistent and stable performance, ensuring a minimum level of effectiveness. Specifically, in Fig. 4 (b), effect-aware scheduling consistently maintains a CPT-based total GoE of $0.29$ or higher throughout the observed period after passing the initial state, i.e., $t=0$. In contrast, using LWGF, WRR, uniform, and Markovian, the CPT-based total GoE could decrease to as low as $0.23$, $0.17$, 0.35-0.35, and 0.35-0.35, respectively, as shown in Fig. 4 (a). This confirms that effect-aware scheduling outperforms the benchmark approaches in consistently maintaining reliable and stable performance.
@@
To elaborate on the computational complexity of the considered model-free DRL algorithms, Fig. 5 illustrates their convergence behavior by representing the number of episodes required to attain stable performance. In this figure, the episodic rewards for each algorithm are smoothed using a $20$-episode moving average to better highlight convergence trends. Under the default settings, model-free PPO demonstrates the fastest convergence, stabilizing approximately $58.33\%$ and $75.66\%$ faster than model-free DQN and A2C, respectively. Notably, DQN and A2C exhibit intermittent performance drops, which can be attributed to their exploration strategies. Nevertheless, these algorithms eventually converge to stable policies as the exploration–exploitation balance is progressively optimized. By applying the dual-loop iterative procedure described in Algorithm 1, obtaining the optimal policies along with the associated Lagrange multiplier requires, on average, $23$ bisection search steps in the outer loop across all model-free algorithms. Specifically, deriving the final optimal policies for query scheduling involves a total of $1,155$ ($21\times 55$), $3,036$ ($23\times 132$), and $5,876$ ($26\times 226$) episodes for PPO, DQN, and A2C, respectively. In contrast, the model-based solution converges in just $378$ ($18\times 21$) iterations, highlighting its significantly lower computational cost. This translates into around $67\%$$93\%$ reduction in the number of training episodes, underscoring the faster convergence and reduced complexity of the model-based alternative.
Figure 5: Convergence performance of different model-free DRL algorithms in terms of episodic rewards and 2020-episode moving average rewards.
To study the interplay between the average CPT-based total GoE and its corresponding reference point, i.e., $\operatorname{GoE}_{\rm ref}$, we plot Fig. 6. This graph depicts the effect of increasing the reference point on the effectiveness and performance of different scheduling approaches. As $\operatorname{GoE}_{\rm ref}$ increases, the effectiveness delivered by all approaches decreases. This decrease can be attributed to the activation of the loss component in the CPT-based value function at higher levels of total GoE. However, as the reference point increases, the performance gap between effect-aware scheduling and the other approaches widens, reaching its peak at $\operatorname{GoE}_{\rm ref}=0.5$. Beyond this critical point, the performance of both model-based and model-free approaches starts to degrade. Notably, after $\operatorname{GoE}_{\rm ref}=0.95$, the benchmark approaches outperform all effect-aware ones.
The reversal likely arises because increasing the reference point compresses the net reward differences across the $\lvert\mathcal{S}\rvert=256$ states, which in turn distorts the optimal value of the Lagrange multiplier. Consequently, this leads to misalignment in the scheduling policies, both model-based and model-free, with respect to the intended effectiveness objectives. This degradation is a direct outcome of the CPT-based evaluation, where even slight changes in perceived gains or losses can substantially affect action preferences. The extent of this distortion depends on the specific effect-aware solution, illustrating a key insight of integrating CPT: higher expectations—reflected by an elevated reference point—can paradoxically lead to lower effectiveness due to overly risk-averse or misdirected decisions.
@@
Regarding the scalability of model-free effect-aware scheduling with a large number of attributes, as depicted in Fig. 8, we analyze how increasing the query limit affects effectiveness. Under the constraint $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 1$ in (III-A), the hub is essentially restricted to sending at most one query during each sampling interval $t$. Fig. 9 illustrates the average CPT-based total GoE over $1,000$ time slots as a function of the query limit, using model-free approaches. Their performance is compared against that of LWGF and uniform scheduling. [^4] Setting the query limit to more than one increases the complexity of the model-free solution and reduces its appeal for exploration in this context, as this requires corresponding modifications to the CMDP model in Section III-B. To plot the corresponding curves, we assume $\lvert\mathcal{M}_{k}\rvert=3,\forall k$ and vary the query limit from $1$ to 33.555Given that update packets are typically small, we assume that multiple attribute updates can be transmitted simultaneously through orthogonal channels or using time-division techniques. The figure shows that increasing the query limit enhances effectiveness across all approaches. Notably, model-free effect-aware scheduling consistently outperforms LWGF on average, particularly when $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 2,\forall t$. Both approaches demonstrate a significant performance advantage over uniform scheduling. When $\mathbf{1}_{N}^{T}\mathbf{A}(t)\mathbf{1}_{M}\leq 3$, the hub can query all attributes simultaneously while still adhering to the cost constraint. At this point, the performance of model-free effect-aware scheduling and LWGF converges, becoming nearly identical.
@@ VII Conclusion
In this paper, we studied the query scheduling problem in end-to-end status update systems operating under a pull-based model. We considered multiple SAs observing a source with various attributes and communicating their observations in the form of updates in response to queries from the hub. The hub is responsible for deciding when to query updates and for constructing a knowledge base to enable AAs to achieve their heterogeneous goals. We introduced the GoE metric and formulated the scheduling problem based on this metric, integrating a risk-averse framework. Specifically, we aimed to maximize the expected discounted sum of the CPT-based total GoE while ensuring compliance with a cost constraint. To solve this problem and design effect-aware scheduling policies, we developed model-based and model-free solutions, supported by their respective iterative algorithms. We then evaluated the performance of these solutions in terms of effectiveness and compared them against benchmark methods. Our results demonstrated that the model-based method could improve the efficiency of sent queries by up to $20\%$ while preserving a level of effectiveness comparable to the best benchmark method, namely LWGF. Conversely, model-free solutions showed the potential to enhance long-term effectiveness by up to $9.2\%$ than the other approaches, albeit at the cost of higher computational complexity. This improvement was particularly pronounced under strict cost constraints, where effect-aware scheduling policies can significantly boost effectiveness and outperform others. Our findings highlighted that the model-free solutions excel in scalability while maintaining high levels of effectiveness, making them particularly appealing for large-scale or dynamic decision-making systems.